given the values of x and y, find the equation of variation in which y varies directly as x. Y =8, when x=5
Answer by stanbon(75887)

given the values of x and y, find the equation of variation in which y varies directly as x. Y =8, when x=5
------------
y = kx
---
solve for "k" given that y = 8 when x = 5
8 = k*5
k = (8/5)
------
Equation:
y = (8/5)x
=================
Cheers,
Stan H.
-----------------